Suppose a function

X

on

R

is given as

X

(

ω

) =

1

5

ω

2

, ω


=

7

1 +

ω

2

, ω >


1. Identify the range of

X

.

2. Find

X

1

(

B

) when

B

= (0

,

1),

B

= [1

,

3],

B

= 4,

B

= (2

,

4)




1
Expert's answer
2021-09-01T06:45:01-0400

Part 1

"\mathrm{The\:set\:of\:values\:of\:the\:dependent\:variable\:for\:which\:a\:function\:is\:defined}\\
\mathrm{For\:a\:parabola}\:ax^2+bx+c\:\mathrm{with\:Vertex}\:\left(x_v,\:y_v\right)\\
\mathrm{If}\:a<0\:\mathrm{the\:range\:is}\:f\left(x\right)\le \:y_v\\
a=-1,\:\mathrm{Vertex}\:\left(x_v,\:y_v\right)=\left(0,\:15\right)\\
f\left(x\right)\le \:15\\
\mathrm{Range\:of\:}15-x^2:\quad \begin{bmatrix}\mathrm{Solution:}\:&\:f\left(x\right)\le \:15\:\\ \:\mathrm{Interval\:Notation:}&\:(-\infty \:,\:15]\end{bmatrix}"


Part 2

The function is "71+w^2dw"

"X=71+w^2\\
\implies \mathrm{A\:function\:g\:is\:the\:inverse\:of\:function\:f\:if\:for}\:y=f\left(x\right),\:\:x=g\left(y\right)\:\\
X=71+w^2\\
w=71+X^2\\
X^{-1}=\sqrt{w-71}"